x= 3.1, 9.4, 1.2, 1.0, 9.0, 5.0, 3.4, 7.4, 0.1, 7.5
y= 8.9, 15.0, 4.8, 6.0, 14.9, 11.9, 9.8, 15.0, 4.7, 13.0

what is the coefficient of determination for this data set?
1
Expert's answer
2020-06-28T17:20:04-0400

Table for Calculating Coefficient of Determination




We have, coefficient of determination = r2 ,where r is the correlation coefficient between x and y.


Now, "r=\frac{n\sum xy-(\sum x)(\sum y)}{\sqrt(n\sum x^2-(\sum x)^2)(n\sum y^2-(\sum y)^2)}"


Here, "n=10,\sum x =47.1, \sum y=104, \sum x^2=328.99, \sum y^2=1239, \sum xy=616.24"


Applying the formula,


"r=\frac{10\times 616.24-47.1\times104}{\sqrt(10\times328.99-(47.1)^2)(10\times1239-(104)^2)}" = 0.9733 (rounded to 4 decimal places)


"\therefore" Coefficient of determination = r2 = (0.9733)2 = 0.9473 (rounded to 4 decimal places)


Answer: The coefficient of determination of the given data is 0.9473.